OpenStudy (donnie1999):

If f(x) = 2x + 3 and g(x) = x^2 + 1, find f(g(3)).

OpenStudy (donnie1999):

@dcskaf

OpenStudy (anonymous):

what did you get?

OpenStudy (donnie1999):

1 sec

OpenStudy (donnie1999):

i got 82?

OpenStudy (anonymous):

what steps did you take?

OpenStudy (anonymous):

f(x) = 2 x+3 and g(x) = x^2+1

OpenStudy (donnie1999):

that the problem i dont know what steps to take?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

look at the second equation f(g(3))

OpenStudy (donnie1999):

would it be 23?

OpenStudy (anonymous):

it says x in g is 3

OpenStudy (anonymous):

and x in f(x) is equal to the result of g(3)

OpenStudy (anonymous):

so first g(3)=3^2+1

OpenStudy (donnie1999):

k

OpenStudy (anonymous):

which is equal to 10 which plug into f(x) f(x)=2(10)+3

OpenStudy (anonymous):

so yes it is 23

OpenStudy (donnie1999):

so would it be 47?

OpenStudy (anonymous):

Donnie watcha need?

OpenStudy (anonymous):

do you get it?

OpenStudy (donnie1999):

@dcskaf do u know if 47 would be wright?

OpenStudy (anonymous):

no it is 23

OpenStudy (donnie1999):

lol srry i did not see that